# DOCTOR: A Multi-Disease Detection Continual Learning Framework Based on Wearable Medical Sensors

### v4

@@ 1 Introduction
Physical illnesses and mental health problems impact the well-being of billions of people around the globe. For centuries, disease detection has been a stressful and time-consuming process for patients. Patients have to visit a physician and undergo a series of physical and even invasive examinations. Fortunately, the emergence of wearable medical sensors (WMSs) and modern advances in artificial intelligence (AI) and machine learning (ML) point to a promising approach to address these problems. WMSs enable continuous monitoring of physiological signals in a passive, user-transparent, and non-invasive manner. A sophisticated ML model can then analyze the data captured by WMSs and perform efficient and effective disease detection. This enables disease detection in an out-of-clinic scenario [11, 41, 14, 29, 9, 2, 27]. However, conventional ML-driven disease detection methods rely on customizing an ML model for each disease and its associated WMS data. They lack adaptability to changes in data distributions and the addition of new classification classes for a given disease. For instance, the ML model trained with physiological signals collected from the elderly might not accurately detect the disease in young adults. The model trained with data from healthy individuals and symptomatic COVID-19 patients might be unable to detect the virus in asymptomatic patients. Moreover, when the model is retrained with new data to adapt to data domain shifts or new classification classes, the model may overfit on new data. Hence, the model may suffer from performance deterioration for previously learned missions. This phenomenon is well known as catastrophic forgetting [26]. In addition, the model trained for detecting one disease cannot be used to detect others. Therefore, new models need to be designed and trained from scratch to detect new diseases. Nevertheless, placing multiple disease-detecting models on edge devices (where WMSs reside) increases memory footprint and drains the battery much faster. Consequently, it makes conventional ML-driven disease detection methods suboptimal.
@@
To address the aforementioned challenges, we propose DOCTOR, a multi-disease detection CL framework based on WMSs. Fig. 1 gives an overview of our framework and compares it to conventional ML-driven disease detection methods. DOCTOR performs multi-disease detection with a deep neural network (DNN) and a replay-style CL algorithm. The CL algorithm mitigates catastrophic forgetting in the tabular data domain by training the DNN in an exemplar- or a generative-replay fashion. It enables the framework to continually learn new missions with a single DNN where different data distributions, classification classes, and disease detection tasks are introduced sequentially. Our framework is applicable to any generic DNN model, including recurrent neural networks (RNNs) and long short-term memory (LSTM) networks. In addition, DOCTOR does not require manual feature extraction procedures. It operates directly on sequential time series tabular data collected from commercially available WMSs.
@@ 4.2.4 Multi-disease Detection Application
The multi-headed architecture allows DOCTOR to generate individual classification probability distribution for each detection head. Therefore, DOCTOR can perform detection for different diseases in parallel. Fig. 4 demonstrates how DOCTOR can be used for multi-disease detection. At test time, physiological signals are collected with commercially available WMSs and sent to the framework. The framework then preprocesses the raw data and performs disease detection with the multi-headed DNN model. Finally, DOCTOR outputs the detection result for each disease from each detection head. Therefore, patients can be informed about which diseases are detected from their physiological signals with a single examination. For example, a patient may learn that he or she is asymptomatic-positive for COVID-19, Type-2 diabetic, and mentally depressed through a single test.
@@ 5.1 Datasets
The DiabDeep dataset contains various data features obtained from 25 non-diabetic individuals, 14 Type-I diabetic patients, and 13 Type-2 diabetic patients. All the data features were recorded as continuous values. The physiological signals of the participants were collected with an Empatica E4 smartwatch. Additional electromechanical and ambient environmental data were recorded with a Samsung Galaxy S4 smartphone. These additional data were proven to provide information that offers diagnostic insights through user habit tracking [40]. For example, they can assist in sensing body movement and calibrating physiological signal collection. Table 2 lists the data features that were collected in this dataset.
@@ 5.3 Implementation Details
DOCTOR exploits a single DNN model and the proposed replay-style CL algorithm to learn various disease detection tasks incrementally in the tabular data domain. Any generic DNN model that performs classification on sequential time series data can be employed in our framework, such as RNNs and LSTM networks. However, these networks require more training data, computational resources, and time for training. Thus, we choose to implement a multilayer perceptron (MLP) model in our framework.
@@ 6 Experimental Results
We evaluate DOCTOR’s efficacy under three different CL settings: domain-incremental, class-incremental, and task-incremental. In addition, we examine the performance of all three replay-style CL methods in DOCTOR: DP, SDG with GMME, and SDG with KDE. We conduct experiments with the CovidDeep, DiabDeep, and MHDeep datasets introduced in Section 5.1. We report the performance of all frameworks with the evaluation metrics mentioned in Section 3.2, including average accuracy (), average F1-score (F1-score), and BWT. In addition, we report the test accuracy () for each learned mission and the required buffer size in megabytes (MB) for all frameworks. Moreover, we record the KS test statistic (described in Section 4.3.2) for the two SDG methods in DOCTOR. Finally, we demonstrate the potential of the multi-disease detection application of DOCTOR and conduct an ablation study.
@@
| Datasets | Frameworks | F1-score | BWT | KS Test Statistic | Buffer Size (MB) | ||||
@@ 6.1 Domain-incremental Continual Learning Scenario
In the domain-incremental scenario, new missions just contain data from different probability distributions but with the same classification classes as the given task. Hence, the MLP models do not need to expand their output layers to accommodate new missions. Table 3 shows the experimental results for all frameworks in the domain-incremental CL scenario. As we can see from the table, for all three datasets, the MLP model suffers from significant performance deterioration for when the naive fine-tuning framework is used, due to catastrophic forgetting. This results in poor , F1-score, and BWT compared to the other frameworks. DOCTOR far outperforms the naive fine-tuning and LwF frameworks, and achieves very competitive , F1-score, and BWT relative to the ideal joint-training framework for all three datasets. Moreover, DOCTOR even outperforms the ideal scenario in some cases due to the generalization gained through learning from synthetic data. Next, we examine the KS test statistics when DOCTOR employs generative replay with the SDG module. When GMME and KDE both yield similar KS test statistics for the CovidDeep and MHDeep datasets, DOCTOR achieves a very similar test accuracy for with the synthetic data generated using each method. However, for the DiabDeep dataset, the KDE method results in a lower KS test statistic and hence a higher test accuracy for . This validates our design decision described in Section 4.3.2 in choosing the estimation method that yields a lower KS test statistic for generative replay.
In summary, DOCTOR achieves a 0.990 , a 0.995 F1-score, and a -0.007 BWT on the CovidDeep dataset, a 0.957 , a 0.961 F1-score, and a -0.004 BWT on the DiabDeep dataset, and a 0.887 , a 0.977 F1-score, and a -0.005 BWT on the MHDeep dataset. This demonstrates that DOCTOR can incrementally adapt to data from new distributions while maintaining the knowledge learned from previous ones. The best DOCTOR results are shown in bold.
@@ 6.2 Class-incremental Continual Learning Scenario
In the class-incremental CL experiments, we assume two missions where contains new classification classes not seen in for the same disease detection task. For CovidDeep, we include only the healthy individuals and symptomatic patients in , whereas only contains the asymptomatic patients. For DiabDeep, includes the healthy individuals and Type-I diabetic patients, whereas comprises solely the Type-2 diabetic patients. For MHDeep, we include the healthy participants and participants with major depressive disorder in , whereas consists of participants with bipolar depressive disorder and schizoaffective disorder. As before, within each mission, we prepare the training, validation, and test sets by taking the first 70%, the next 10%, and the last 20% of the sequential time series data instances from each patient with no time overlap. Similarly, we apply SMOTE to the training sets in both missions to address the data imbalance issue.
@@
Table 4 presents the experimental results for all frameworks in the class-incremental CL scenario. As can again be seen from the table, due to catastrophic forgetting, the MLP model suffers from significant performance degradation on when the naive fine-tuning framework is used in all three datasets. However, DOCTOR significantly outperforms the naive fine-tuning and LwF frameworks. Moreover, due to fine-tuning with the most informative preserved data and the generalization gained through learning from synthetic data, it even achieves higher and BWT and a competitive F1-score relative to the ideal joint-training framework for all three datasets. On the other hand, as shown in the table, the estimation method that yields a lower KS test statistic results in a better test accuracy for for all datasets. This again validates the design decision we made regarding the KS test statistic in Section 4.3.2.
| Datasets | Frameworks | F1-score | BWT | KS Test Statistic | Buffer Size (MB) | ||||
@@
To summarize, DOCTOR achieves a 0.986 , a 0.991 F1-score, and a -0.008 BWT on the CovidDeep dataset, a 0.921 , a 0.929 F1-score, and a -0.025 BWT on the DiabDeep dataset, and a 0.895 , a 0.983 F1-score, and a -0.001 BWT on the MHDeep dataset. This demonstrates the efficacy of DOCTOR in learning new classification classes while preserving prior knowledge. Again, the best DOCTOR results are shown in bold.
@@ 6.3 Task-incremental Continual Learning
| Datasets | Frameworks | F1-score | BWT | KS Test Statistic | Buffer Size (MB) | |||||
@@
DOCTOR achieves a 0.962 , a 0.964 F1-score, and a 0.002 BWT after continually learning two disease detection tasks, and a 0.928 , a 0.969 F1-score, and a 0.002 BWT after consecutively learning to detect three diseases. This demonstrates DOCTOR’s efficacy in the task-incremental scenario. The best DOCTOR results are again shown in bold.
@@ 6.5 Cost Analysis
| CL | Datasets | Subset | F1-score | BWT | Buffer Size | |
@@
6.6 Ablation Study
Next, we present results for several ablation studies.
6.6.1 Informative Subset for Data Preservation
Table 7 shows the experimental results of the ablation study. The best results are shown in bold. We can see that preserving the real training data subset whose average training loss values are above the 50th percentile (the top 50%) results in the best performance (, F1-score, and BWT) in all the experiments. These results demonstrate that training data with higher average training loss values are more informative. Hence, they can better assist the model in retaining the learned knowledge. Therefore, in our DP method, we preserve the real training data subset with the highest average training loss values.
@@ 6.6.2 Amount of Data Preserved
Here, we perform an ablation study to examine the performance (, F1-score, BWT, and buffer size) of the DP CL algorithm versus the amount of preserved real training data. We repeat the CL experiments again as in Section 6.6.1 with only the DP method. However, this time, we repeat the experiments nine times while sweeping the threshold value from the 90th percentile to the 10th percentile in steps of 10 to preserve the top 10%, top 20%, , and top 90% of the real training data.
Fig. 8, 9, and 10 show the experimental results of the ablation study in domain-, class-, and task-incremental CL experiments, respectively. This ablation study aims to find the optimal trade-off between the buffer size needed and other performance metrics. As we can see from the figures, preserving the top 30% of real training data leads to minimum buffer requirements with acceptable performance loss. In general, preserving more real training data for replay should give better performance in CL. However, this does not always hold true. As the figures show, there might be a drop in , F1-score, and BWT when we preserve more real training data. This is because the framework runs the risk of overfitting to previous missions and failing to learn the new ones when more preserved data are replayed. Moreover, preserving more data requires more memory storage and makes the framework harder to scale to many consecutive missions. Therefore, we set the threshold value to the 70th percentile to preserve the top 30% of the real training data in the DP method.
@@ 6.6.3 Amount of Synthetic Data
Finally, we report results for an ablation study of the performance (, F1-score, and BWT) of the SDG CL algorithm versus the amount of synthetic data used. Since the SDG module generates synthetic data on the fly with the stored probability density estimation models during training, no extra memory buffer is needed. We repeat the experiments described in Sections 6.1, 6.2, and 6.3 with only the SDG method under the same settings for all three datasets. For the amount of synthetic data generated, we target 20%, 40%, , and 100% of the amount of real training data.

### v5

@@ 1 Introduction
Physical illnesses and mental health problems impact the well-being of billions of people around the globe. For centuries, disease detection has been a stressful and time-consuming process for patients. Patients have to visit a physician and undergo a series of physical and even invasive examinations. Fortunately, the emergence of wearable medical sensors (WMSs) and modern advances in artificial intelligence (AI) and machine learning (ML) point to a promising approach to address these problems. WMSs enable continuous monitoring of physiological signals in a passive, user-transparent, and non-invasive manner. A sophisticated ML model can then analyze the data captured by WMSs and perform efficient and effective disease detection. This enables disease detection in an out-of-clinic setting [11, 43, 14, 29, 9, 2, 27].
In real-world scenarios, health record datasets are governed by numerous laws and restrictions to ensure patient data privacy. Hence, smart healthcare application designers are often granted access to these datasets for only a limited period of time with restricted or no permission to preserve data from them. Conventional ML-driven disease detection methods rely on customizing an ML model for each disease based on its associated WMS dataset. Once new datasets collected from different domains, containing new classification classes, or gathered for different disease detection tasks become available, conventional methods require access to both new and old datasets to update the models. Simply retraining the models with new datasets with conventional methods causes the models to overfit on new data and suffer from performance deterioration for previously learned missions. This phenomenon is well known as catastrophic forgetting [26]. Therefore, losing access to previously used datasets hinders conventional methods from adapting to changes in data distributions and additions of new classification classes for a given disease. In addition, it reduces the flexibility of a trained model for learning to detect a distinct disease. For instance, an ML model trained with a dataset collected from the elderly might not accurately detect the disease in young adults. A model trained with data gathered from healthy individuals and symptomatic COVID-19 patients might be unable to detect the virus in asymptomatic patients. These models suffer from performance deterioration on previous datasets when they are retained using conventional methods on new datasets after losing access to prior ones. In addition, with conventional methods, a model trained for detecting one disease cannot be updated to detect other diseases while maintaining performance for the original disease after losing access to its dataset. Therefore, new models need to be designed and trained from scratch to detect new diseases. However, placing multiple disease-detecting models on edge devices (where WMSs reside) significantly increases memory footprint and drains the battery much faster. Consequently, it makes conventional ML-driven disease detection methods suboptimal.
@@
To address the aforementioned challenges, we propose DOCTOR, a multi-disease detection CL framework based on WMSs. Fig. 1 gives an overview of our framework and compares it to conventional ML-driven disease detection methods. DOCTOR performs multi-disease detection with a deep neural network (DNN) and a replay-style CL algorithm. The CL algorithm mitigates catastrophic forgetting in the tabular data domain by training the DNN in an exemplar- or a generative-replay fashion. It enables the framework to continually learn new missions with a single DNN where different data distributions, classification classes, and disease detection tasks are introduced sequentially. Our framework is applicable to other generic DNN models, including recurrent neural networks (RNNs) [34] and long short-term memory (LSTM) networks [33]. In addition, DOCTOR does not require manual feature extraction procedures. It operates directly on sequential time series tabular data collected from commercially available WMSs.
@@ 4.2.4 Multi-disease Detection Application
The multi-headed architecture allows DOCTOR to generate individual classification probability distribution for each detection head. Therefore, DOCTOR can perform detection for different diseases in parallel. Fig. 4 demonstrates how DOCTOR can be used for multi-disease detection. At test time, physiological signals are collected with commercially available WMSs and sent to the framework. The framework then preprocesses the raw data and performs disease detection with the multi-headed DNN model. Finally, DOCTOR outputs the detection result for each disease from each detection head. Therefore, patients can be informed about which diseases are detected from their physiological signals with a single examination. For example, a patient may learn that he or she is asymptomatic-positive for COVID-19, Type-II diabetic, and mentally depressed through a single test.
@@ 5.1 Datasets
The DiabDeep dataset contains various data features obtained from 25 non-diabetic individuals, 14 Type-I diabetic patients, and 13 Type-II diabetic patients. All the data features were recorded as continuous values. The physiological signals of the participants were collected with an Empatica E4 smartwatch. Additional electromechanical and ambient environmental data were recorded with a Samsung Galaxy S4 smartphone. These additional data were proven to provide information that offers diagnostic insights through user habit tracking [42]. For example, they can assist in sensing body movement and calibrating physiological signal collection. Table 2 lists the data features that were collected in this dataset.
@@ 5.3 Implementation Details
DOCTOR exploits a single DNN model and the proposed replay-style CL algorithm to learn various disease detection tasks incrementally in the tabular data domain. Other generic DNN models that perform classification on sequential time series data can be employed in our framework, such as RNNs and LSTM networks (see an ablation study in Section 6.6.4). However, these networks generally require more training data, computational resources, and time for training. Thus, we choose to implement a multilayer perceptron (MLP) model in our framework.
@@ 6 Experimental Results
We evaluate DOCTOR’s efficacy under three different CL settings: domain-incremental, class-incremental, and task-incremental. In addition, we examine the performance of all three replay-style CL methods in DOCTOR: DP, SDG with GMME, and SDG with KDE. We conduct experiments with the CovidDeep, DiabDeep, and MHDeep datasets introduced in Section 5.1. We report the performance of all frameworks with the evaluation metrics mentioned in Section 3.2, including average accuracy (), average F1-score (F1-scoreavg), and BWT. In addition, we report the test accuracy () for each learned mission and the required buffer size in megabytes (MB) for all frameworks. Moreover, we record the KS test statistic (described in Section 4.3.2) for the two SDG methods in DOCTOR. Finally, we demonstrate the potential of the multi-disease detection application of DOCTOR and conduct an ablation study.
@@
| Datasets | Frameworks | F1-scoreavg | BWT | KS Test Statistic | Buffer Size (MB) | ||||
@@ 6.1 Domain-incremental Continual Learning Scenario
In the domain-incremental scenario, new missions just contain data from different probability distributions but with the same classification classes as the given task. Hence, the MLP models do not need to expand their output layers to accommodate new missions. Table 3 shows the experimental results for all frameworks in the domain-incremental CL scenario. As we can see from the table, for all three datasets, the MLP model suffers from significant performance deterioration for when the naive fine-tuning framework is used, due to catastrophic forgetting. This results in poor , F1-scoreavg, and BWT compared to the other frameworks. DOCTOR far outperforms the naive fine-tuning and LwF frameworks, and it achieves very competitive , F1-scoreavg, and BWT relative to the ideal joint-training framework for all three datasets. Moreover, DOCTOR even outperforms the ideal scenario in some cases due to the generalization gained through learning from synthetic data. Next, we examine the KS test statistics when DOCTOR employs generative replay with the SDG module. When GMME and KDE both yield similar KS test statistics for the CovidDeep and MHDeep datasets, DOCTOR achieves a very similar test accuracy for with the synthetic data generated using each method. However, for the DiabDeep dataset, the KDE method results in a lower KS test statistic and hence a higher test accuracy for . This validates our design decision described in Section 4.3.2 in choosing the estimation method that yields a lower KS test statistic for generative replay.
In summary, DOCTOR achieves a 0.990 , a 0.995 F1-scoreavg, and a -0.007 BWT on the CovidDeep dataset, a 0.957 , a 0.961 F1-scoreavg, and a -0.004 BWT on the DiabDeep dataset, and a 0.887 , a 0.977 F1-scoreavg, and a -0.005 BWT on the MHDeep dataset. This demonstrates that DOCTOR can incrementally adapt to data from new distributions while maintaining the knowledge learned from previous ones. The best DOCTOR results are shown in bold.
@@ 6.2 Class-incremental Continual Learning Scenario
In the class-incremental CL experiments, we assume two missions where contains new classification classes not seen in for the same disease detection task. For CovidDeep, we include only the healthy individuals and symptomatic patients in , whereas only contains the asymptomatic patients. For DiabDeep, includes the healthy individuals and Type-I diabetic patients, whereas comprises solely the Type-II diabetic patients. For MHDeep, we include the healthy participants and participants with major depressive disorder in , whereas consists of participants with bipolar depressive disorder and schizoaffective disorder. As before, within each mission, we prepare the training, validation, and test sets by taking the first 70%, the next 10%, and the last 20% of the sequential time series data instances from each patient with no time overlap. Similarly, we apply SMOTE to the training sets in both missions to address the data imbalance issue.
@@
Table 4 presents the experimental results for all frameworks in the class-incremental CL scenario. As can again be seen from the table, due to catastrophic forgetting, the MLP model suffers from significant performance degradation on when the naive fine-tuning framework is used in all three datasets. However, DOCTOR significantly outperforms the naive fine-tuning and LwF frameworks. Moreover, due to fine-tuning with the most informative preserved data and the generalization gained through learning from synthetic data, it even achieves higher and BWT and a competitive F1-scoreavg relative to the ideal joint-training framework for all three datasets. On the other hand, as shown in the table, the estimation method that yields a lower KS test statistic results in a better test accuracy for for all datasets. This again validates the design decision we made regarding the KS test statistic in Section 4.3.2.
| Datasets | Frameworks | F1-scoreavg | BWT | KS Test Statistic | Buffer Size (MB) | ||||
@@
To summarize, DOCTOR achieves a 0.986 , a 0.991 F1-scoreavg, and a -0.008 BWT on the CovidDeep dataset, a 0.921 , a 0.929 F1-scoreavg, and a -0.025 BWT on the DiabDeep dataset, and a 0.895 , a 0.983 F1-scoreavg, and a -0.001 BWT on the MHDeep dataset. This demonstrates the efficacy of DOCTOR in learning new classification classes while preserving prior knowledge. Again, the best DOCTOR results are shown in bold.
@@ 6.3 Task-incremental Continual Learning
| Datasets | Frameworks | F1-scoreavg | BWT | KS Test Statistic | Buffer Size (MB) | |||||
@@
DOCTOR achieves a 0.962 , a 0.964 F1-scoreavg, and a 0.002 BWT after continually learning two disease detection tasks, and a 0.928 , a 0.969 F1-scoreavg, and a 0.002 BWT after consecutively learning to detect three diseases. This demonstrates DOCTOR’s efficacy in the task-incremental scenario. The best DOCTOR results are again shown in bold.
@@ 6.5 Cost Analysis
6.6 Ablation Study
Next, we present results for several ablation studies.
6.6.1 Informative Subset for Data Preservation
Table 7 shows the experimental results of the ablation study. The best results are shown in bold. We can see that preserving the real training data subset whose average training loss values are above the 50th percentile (the top 50%) results in the best performance (, F1-scoreavg, and BWT) in all the experiments. These results demonstrate that training data with higher average training loss values are more informative. Hence, they can better assist the model in retaining the learned knowledge. Therefore, in our DP method, we preserve the real training data subset with the highest average training loss values.
| CL | Datasets | Subset | F1-scoreavg | BWT | Buffer Size | |
@@ 6.6.2 Amount of Data Preserved
Here, we perform an ablation study to examine the performance (, F1-scoreavg, BWT, and buffer size) of the DP CL algorithm versus the amount of preserved real training data. We repeat the CL experiments again as in Section 6.6.1 with only the DP method. However, this time, we repeat the experiments nine times while sweeping the threshold value from the 90th percentile to the 10th percentile in steps of 10 to preserve the top 10%, top 20%, , and top 90% of the real training data.
Fig. 8, 9, and 10 show the experimental results of the ablation study in domain-, class-, and task-incremental CL experiments, respectively. This ablation study aims to find the optimal trade-off between the buffer size needed and other performance metrics. As we can see from the figures, preserving the top 30% of real training data leads to minimum buffer requirements with acceptable performance loss. In general, preserving more real training data for replay should give better performance in CL. However, this does not always hold true. As the figures show, there might be a drop in , F1-scoreavg, and BWT when we preserve more real training data. This is because the framework runs the risk of overfitting to previous missions and failing to learn the new ones when more preserved data are replayed. Moreover, preserving more data requires more memory storage and makes the framework harder to scale to many consecutive missions. Therefore, we set the threshold value to the 70th percentile to preserve the top 30% of the real training data in the DP method.
@@ 6.6.3 Amount of Synthetic Data
Here, we report results for an ablation study of the performance (, F1-scoreavg, and BWT) of the SDG CL algorithm versus the amount of synthetic data used. Since the SDG module generates synthetic data on the fly with the stored probability density estimation models during training, no extra memory buffer is needed. We repeat the experiments described in Sections 6.1, 6.2, and 6.3 with only the SDG method under the same settings for all three datasets. For the amount of synthetic data generated, we target 20%, 40%, , and 100% of the amount of real training data.
@@
6.6.4 Alternative Machine Learning Model
Finally, we perform an ablation study to demonstrate DOCTOR’s applicability to other generic DNN models. We repeat the same domain-, class-, and task-incremental CL experiments described in Sections 6.1, 6.2, and 6.3, respectively, with an LSTM network. We adopt the same datasets and preprocessing processes as described in Sections 5.1 and 5.2. For each CL experiment, we follow the same procedures to prepare the training, validation, and test sets as mentioned in Sections 6.1, 6.2, and 6.3. We perform a grid search over various hyperparameters on the validation sets of the three datasets to design the LSTM network. We find that a three-layer LSTM network performs the best in general. Fig. 14 shows the architecture of the final LSTM model used in this ablation study. The first layer is a basic LSTM cell [33] with a hidden state size of 128. Then, it is followed by a fully-connected layer with 128 neurons and ReLU activation. Lastly, it incorporates a multi-headed softmax output layer. We implement this LSTM network in PyTorch for the baseline, naive fine-tuning, LwF, DOCTOR, and ideal joint-training frameworks in all CL scenarios for comparison. The LSTM network in our DOCTOR framework has 163,722 parameters and a size of 625KB. We apply the same training recipe to train all frameworks and adopt the same parameter settings for our DOCTOR framework as described in Section 5.3. Similarly, we perform the experiments on an NVIDIA A100 GPU with CUDA and cuDNN libraries to accelerate the experiments.
Table 8, 9, and 10 show the experimental results for this ablation study in domain-, class-, and task-incremental CL scenarios, respectively. The best DOCTOR results are shown in bold. For all datasets in all three CL scenarios, the naive fine-tuning framework still suffers from significant performance deterioration for previous missions due to catastrophic forgetting. On the other hand, despite using a different DNN, DOCTOR still greatly outperforms the naive fine-tuning and LwF frameworks in all experiments. Moreover, it achieves very competitive performance relative to the ideal joint-training framework in all CL scenarios. In fact, DOCTOR with the LSTM network even performs better than DOCTOR with the MLP model in the domain-incremental CL scenario, but at the cost of a larger model size. This ablation study demonstrates that DOCTOR is directly applicable to other DNN models and can still achieve high disease classification accuracy in various CL scenarios.
| Datasets | Frameworks | F1-scoreavg | BWT | KS Test Statistic | Buffer Size (MB) | ||||
| CovidDeep | Ideal Joint-training | 0.982 | 0.983 | 0.983 | 0.995 | -0.002 | - | 5.8 | |
| DOCTOR | - DP | 0.929 | 0.983 | 0.956 | 0.977 | -0.055 | - | 1.7 | |
| - SDG with GMME | 0.938 | 0.996 | 0.967 | 0.980 | -0.045 | 0.083 | 0 | ||
| - SDG with KDE | 0.978 | 0.999 | 0.988 | 0.994 | -0.006 | 0.079 | 0 | ||
| LwF | 0.897 | 0.999 | 0.948 | 0.964 | -0.085 | - | 0 | ||
| Naive Fine-tuning | 0.786 | 1.000 | 0.893 | 0.936 | -0.198 | - | 0 | ||
| Baseline | 0.983 | 0.819 | 0.901 | 0.937 | - | - | 0 | ||
| DiabDeep | Ideal Joint-training | 0.917 | 0.996 | 0.957 | 0.957 | 0.011 | - | 11.5 | |
| DOCTOR | - DP | 0.915 | 0.996 | 0.956 | 0.956 | 0.008 | - | 3.4 | |
| - SDG with GMME | 0.896 | 0.957 | 0.927 | 0.928 | -0.011 | 0.215 | 0 | ||
| - SDG with KDE | 0.932 | 0.999 | 0.965 | 0.967 | 0.024 | 0.050 | 0 | ||
| LwF | 0.810 | 0.989 | 0.900 | 0.905 | -0.131 | - | 0 | ||
| Naive Fine-tuning | 0.631 | 1.000 | 0.816 | 0.825 | -0.276 | - | 0 | ||
| Baseline | 0.907 | 0.372 | 0.640 | 0.649 | - | - | 0 | ||
| MHDeep | Ideal Joint-training | 0.831 | 0.953 | 0.892 | 0.974 | -0.012 | - | 13.6 | |
| DOCTOR | - DP | 0.820 | 0.955 | 0.888 | 0.977 | -0.023 | - | 4.1 | |
| - SDG with GMME | 0.832 | 0.898 | 0.865 | 0.967 | -0.011 | 0.041 | 0 | ||
| - SDG with KDE | 0.832 | 0.924 | 0.878 | 0.967 | -0.011 | 0.036 | 0 | ||
| LwF | 0.657 | 0.943 | 0.800 | 0.883 | -0.197 | - | 0 | ||
| Naive Fine-tuning | 0.420 | 0.983 | 0.702 | 0.825 | -0.423 | - | 0 | ||
| Baseline | 0.843 | 0.114 | 0.479 | 0.755 | - | - | 0 | ||
| Datasets | Frameworks | F1-scoreavg | BWT | KS Test Statistic | Buffer Size (MB) | ||||
| CovidDeep | Ideal Joint-training | 0.978 | 0.992 | 0.985 | 0.990 | -0.020 | - | 5.0 | |
| DOCTOR | - DP | 0.807 | 0.983 | 0.895 | 0.914 | -0.192 | - | 1.5 | |
| - SDG with GMME | 0.961 | 0.984 | 0.973 | 0.981 | -0.037 | 0.166 | 0 | ||
| - SDG with KDE | 0.987 | 0.972 | 0.980 | 0.992 | -0.011 | 0.081 | 0 | ||
| LwF | 0.107 | 1.000 | 0.553 | 0.837 | -0.892 | - | 0 | ||
| Naive Fine-tuning | 0 | 1.000 | 0.500 | 0.837 | -0.998 | - | 0 | ||
| Baseline | 0.998 | 0 | 0.499 | 0.766 | - | - | 0 | ||
| DiabDeep | Ideal Joint-training | 0.882 | 0.943 | 0.913 | 0.922 | -0.042 | - | 9.2 | |
| DOCTOR | - DP | 0.896 | 0.970 | 0.933 | 0.945 | -0.023 | - | 2.8 | |
| - SDG with GMME | 0.924 | 0.912 | 0.918 | 0.921 | 0.005 | 0.063 | 0 | ||
| - SDG with KDE | 0.919 | 0.970 | 0.944 | 0.953 | 0 | 0.211 | 0 | ||
| LwF | 0 | 1.000 | 0.500 | 0.647 | -0.958 | - | 0 | ||
| Naive Fine-tuning | 0 | 1.000 | 0.500 | 0.647 | -0.919 | - | 0 | ||
| Baseline | 0.919 | 0 | 0.459 | 0.696 | - | - | 0 | ||
| MHDeep | Ideal Joint-training | 0.892 | 0.874 | 0.883 | 0.983 | -0.029 | - | 8.3 | |
| DOCTOR | - DP | 0.893 | 0.889 | 0.891 | 0.980 | -0.045 | - | 2.5 | |
| - SDG with GMME | 0.927 | 0.867 | 0.897 | 0.953 | -0.012 | 0.042 | 0 | ||
| - SDG with KDE | 0.907 | 0.889 | 0.898 | 0.981 | -0.031 | 0.209 | 0 | ||
| LwF | 0.242 | 0.849 | 0.545 | 0.772 | -0.645 | - | 0 | ||
| Naive Fine-tuning | 0 | 0.904 | 0.452 | 0.775 | -0.938 | - | 0 | ||
| Baseline | 0.938 | 0 | 0.469 | 0.500 | - | - | 0 | ||
| Datasets | Frameworks | F1-scoreavg | BWT | KS Test Statistic | Buffer Size (MB) | |||||
| MHDeep | Ideal Joint-training | 0.993 | 0.920 | 0.838 | 0.917 | 0.965 | 0.014 | - | 21.1 | |
| DOCTOR | - DP | 0.973 | 0.935 | 0.868 | 0.925 | 0.968 | 0.009 | - | 6.3 | |
| DiabDeep | - SDG with GMME | 0.970 | 0.918 | 0.861 | 0.916 | 0.959 | -0.007 | 0.036 | 0 | |
| - SDG with KDE | 0.963 | 0.921 | 0.868 | 0.917 | 0.962 | -0.004 | 0.041 | 0 | ||
| LwF | 0.307 | 0.921 | 0.877 | 0.702 | 0.916 | -0.349 | - | 0 | ||
| CovidDeep | Naive Fine-tuning | 0.879 | 0.575 | 0.862 | 0.772 | 0.847 | -0.220 | - | 0 | |
| DiabDeep | Ideal Joint-training | 0.992 | 0.899 | - | 0.946 | 0.949 | 0.007 | - | 7.3 | |
| DOCTOR | - DP | 0.960 | 0.917 | - | 0.938 | 0.951 | -0.012 | - | 2.2 | |
| - SDG with GMME | 0.972 | 0.930 | - | 0.951 | 0.957 | 0 | 0.081 | 0 | ||
| - SDG with KDE | 0.966 | 0.919 | - | 0.943 | 0.952 | -0.006 | 0.079 | 0 | ||
| CovidDeep | LwF | 0.307 | 0.948 | - | 0.628 | 0.892 | -0.671 | - | 0 | |
| Naive Fine-tuning | 0.931 | 0.923 | - | 0.927 | 0.945 | -0.041 | - | 0 | ||
| CovidDeep | Baseline | 0.972 | - | - | 0.972 | 0.985 | - | - | 0 | |